# Mohana Prasad Sathya Moorthy, Jeff Nichols, Yinfei Yang, Zhe Gan Apple

### v1

@@ Acknowledgements
The authors would like to thank Forrest Huang, Zhen Yang, Haoxuan You, Amanda Swearngin and Alexander Toshev for valuable guidance, suggestions, and feedback.
@@ Appendix A Detailed Results on Elementary and Advanced Tasks
Table 6 shows the performance on 9 subtasks $\times$ 5 platforms of the Ferret-UI 2 model. We emprically find that OCR, widget listing and comprehensive description are relatively difficult tasks compared to other tasks within the same task type.
@@
Appendix B Detailed GPT-4o Prompts for Generating Advanced Task Data
In this section, we elaborate how we prompt GPT-4o to generate training data for advanced tasks. In particular, we have the following requirements when generating each type of advanced tasks.
@@ Appendix B Detailed GPT-4o Prompts for Generating Advanced Task Data
      Predict current step for potential navigation paths based on currently visible widgets like tabs, menus and breadcrumbs controls.
Appendix C Advanced Task Examples
@@
Appendix D Additional Inference Examples
Figures 6, 7, 8 and 9 show more qualitative inference examples of Ferret-UI 2 on different platforms. Additionally, in Figure 10, we show an example of Ferret-UI 2 performing multi-step interactions on real-time webpages following GUIDE-style QAs.
Appendix E Resolution Statistics
@@
Appendix F Label Statistics and Mapping Results of Original Data across Platforms
We demonstrate the label statistics of original data across platforms and their mapping results into a uniform label space for better joint training as shown in Figures 11, 12, 13 and 14. The mapping results are obtained via GPT-4 suggestions and additional human reviews. Note that “Other” label after mapping indicates the widget information will be deprecated.

### v2

@@ 5 Conclusions
Ethics Statement
The use of UI understanding data for this paper was approved by our organization’s legal team. We have strictly adhered to ethical standards, ensuring that no private or sensitive information has been used or compromised.
@@ Acknowledgements
The authors would like to thank Forrest Huang, Zhen Yang, Haoxuan You, Amanda Swearngin, Alexander Toshev and Yang Zhao for valuable guidance, suggestions, and feedback.
@@ Appendix A Detailed Results on Elementary and Advanced Tasks
Table 6 shows the performance on 9 subtasks $\times$ 5 platforms of the Ferret-UI 2 model. We empirically find that OCR, widget listing, and comprehensive description are relatively difficult tasks compared to other tasks within the same task type.
@@
Appendix B Grounding Performance on ScreenSpot Benchmark
We evaluate Ferret-UI 2 on the ScreenSpot (Cheng et al., 2024) benchmark for grounding performance on unseen data. Results are shown in Table 7. Ferret-UI 2 achieves 54.0% average accuracy, outperforming CogAgent (Hong et al., 2023) and SeeClick (Cheng et al., 2024), and in particular, achieving good performance on our supported Mobile and Web platforms while achieving fair performance on the unseen Desktop platform.
Table 7: Performance of Ferret-UI 2 on the ScreenSpot benchmark (Cheng et al., 2024). I/W refers to Icon/Widget data type. Note that the data distribution of ScreenSpot is unseen during Ferret-UI 2 training. The results indicate that Ferret-UI 2 achieve superior grounding accuracy on Mobile and Web screenshots, while underperforms other two models on the unseen Desktop screenshots, which means Ferret-UI 2 transfers well to unseen data in supported platforms.
Model Model Size Mobile Desktop Web Avg
Text I/W Text I/W Text I/W
CogAgent 8B 67.0% 24.0% 74.2% 20.0% 70.4% 28.6% 47.4%
SeeClick 9.6B 78.0% 52.0% 72.2% 30.0% 55.7% 32.5% 53.4%
Ferret-UI 2 8B 80.3% 55.4% 52.1% 21.7% 81.2% 33.5% 54.0%
Appendix C Detailed GPT-4o Prompts for Generating Advanced Task Data
In this section, we elaborate on how we prompt GPT-4o to generate training data for advanced tasks. In particular, we have the following requirements when generating each type of advanced task.
@@ Appendix C Detailed GPT-4o Prompts for Generating Advanced Task Data
      Predict the current step for potential navigation paths based on currently visible widgets like tabs, menus, and breadcrumbs controls.
Table 8: Statictics of different multi-platform GUI understanding datasets.
Dataset # Sample Platforms Tasks
OmniAct 9,802 Desktop, Web Code Generation
OS-World 369 Desktop, Web General Control
AITW 715,142 Android(Apps+Web) Navigation
GUI-world 12,379 iPhone, Android, Web, XR, Desktop GUI understanding, instruction following
Ferret-UI 123,702 iPhone, Android
Elementary: Referring, Grounding
Advanced: Function/detailed description,
perception and interaction QA
Ferret-UI 2 529,411 iPhone, Android, iPad, AppleTV, Web
Elementary: Referring, Grounding
Advanced: Comprehensive description,
perception and interaction QA
Appendix D Advanced Task Examples
@@
Appendix E Dataset Comparison
Figure 6: Illustration of the data generation pipeline of Ferret-UI (You et al., 2024).
In this section, we compare the dataset of Ferret-UI 2 with previous multi-platform datasets as shown in Table 8. In particular, we outline the data generation pipeline of Ferret-UI (You et al., 2024) in Figure 6. Through a comparison of Figure 2 and 6, we highlight the following key differences between Ferret-UI 2 and Ferret-UI:
  • Multi-platform support: The training data of Ferret-UI 2 consists of 5 platforms compared to 2 platforms of Ferret-UI.
  • Raw annotation qualities: The majority part of bounding boxes, labels, and on-screen text annotations of Ferret-UI 2 are either extracted from source data or annotated by humans, while these annotations of Ferret-UI are all generated from model detection.
  • Bounding box prompting: When generating advanced tasks, Ferret-UI 2 uses GPT-4o+SoM visual prompting for bounding boxes, while Ferret-UI uses purely textual prompting containing coordinates, bbox labels, and text annotations.
  • Advanced task quality: Due to the constraint of textual prompting, the Ferret-UI pipeline cannot perceive the visual content of UI elements, thus limiting the quality and content diversity of generated advanced tasks.
Appendix F Resolution Statistics
@@
Appendix G Label Statistics and Mapping Results of Original Data across Platforms
We demonstrate the label statistics of original data across platforms and their mapping results into a uniform label space for better joint training as shown in Figures 7, 8, 9 and 10. The mapping results are obtained via GPT-4 suggestions and additional human reviews. Note that the “Other” label after mapping indicates the widget information will be deprecated.
@@ Appendix G Label Statistics and Mapping Results of Original Data across Platforms
Appendix H Additional Inference Examples
In Figure 11 to 14, we show additional qualitative inference examples of Ferret-UI 2 on different platforms. Additionally, in Figure 15, we show an example of Ferret-UI 2 performing multi-step interactions on real-time webpages following GUIDE-style QAs.